# The Master Key Filters Hypothesis: Deep Filters Are General in DS-CNNs

### v1

@@ Introduction
Results. The results of these experiments are presented in Table 6. Transferring filters from the larger variants of ConvNeXt trained on ImageNet leads to better accuracy improvements compared to the small femto variant. Interestingly, the transferred filters from HorNet perform exceptionally well, despite being from a completely different model architecture and size.
@@
Our experiments, spanning semantically divided ImageNet, cross-domain, and cross-architecture transfers, reveal that these filters maintain their generality even in deeper layers, challenging prevailing notions from traditional CNNs, that there is a transition from general to specialized filters, and filters get increasingly specialized in deeper layers of the network.
When transferring the pointwise layers, we observed convergence issues across all datasets, even in selffer models, suggesting optimizer difficulties when these layers are frozen. This may be attributed to the higher parameter count in pointwise layers, potential sparsity effects, and restricted permutation symmetries compared to depthwise layers. These findings align with prior work on fragile co-adaptation in neural networks (Yosinski et al. 2014), where freezing certain layers can create challenging loss landscapes for training the remaining parameters. Future work could further investigate the specific mechanisms behind these optimization challenges.
@@ Acknowledgments
Checklist
Does this paper rely on one or more datasets? (yes)
A motivation is given for why the experiments are conducted on the selected datasets (yes)
All novel datasets introduced in this paper are included in a data appendix. (NA)
All novel datasets introduced in this paper will be made publicly available upon publication of the paper with a license that allows free usage for research purposes. (NA)
All datasets drawn from the existing literature (potentially including authors’ own previously published work) are accompanied by appropriate citations. (yes)
All datasets drawn from the existing literature (potentially including authors’ own previously published work) are publicly available. (yes)
All datasets that are not publicly available are described in detail, with explanation why publicly available alternatives are not scientifically satisficing. (NA)
Does this paper include computational experiments? (yes) If yes, please complete the list below.
Any code required for pre-processing data is included in the appendix. (yes).
All source code required for conducting and analyzing the experiments is included in a code appendix. (yes)
All source code required for conducting and analyzing the experiments will be made publicly available upon publication of the paper with a license that allows free usage for research purposes. (yes)
All source code implementing new methods have comments detailing the implementation, with references to the paper where each step comes from (yes)
If an algorithm depends on randomness, then the method used for setting seeds is described in a way sufficient to allow replication of results. (NA)
This paper specifies the computing infrastructure used for running experiments (hardware and software), including GPU/CPU models; amount of memory; operating system; names and versions of relevant software libraries and frameworks. (yes)
This paper formally describes evaluation metrics used and explains the motivation for choosing these metrics. (yes)
This paper states the number of algorithm runs used to compute each reported result. (yes)
Analysis of experiments goes beyond single-dimensional summaries of performance (e.g., average; median) to include measures of variation, confidence, or other distributional information. (no)
The significance of any improvement or decrease in performance is judged using appropriate statistical tests (e.g., Wilcoxon signed-rank). (no)
This paper lists all final (hyper-)parameters used for each model/algorithm in the paper’s experiments. (yes)
This paper states the number and range of values tried per (hyper-) parameter during development of the paper, along with the criterion used for selecting the final parameter setting. (NA)

### v2

@@ Introduction
To further challenge the notion of layer specialization, we conducted the reverse of Yosinski’s experiment—freezing layers from the end to n while training earlier layers in Figure 4. Following the same interpretation as Yosinski’s framework, these results would paradoxically suggest that later layers are general while initial layers are specialized.
@@
Results. The results of these experiments are presented in Table 6. Transferring filters from larger variants of ConvNeXt trained on ImageNet leads to better accuracy improvements compared to the smaller variant. Interestingly, the transferred filters from HorNet perform exceptionally well, despite being from a completely different model architecture and size.
@@
Our experiments, spanning semantically divided ImageNet, cross-domain, and cross-architecture transfers, reveal that these filters maintain generality even in deeper layers, challenging prevailing notions that there is a transition from general to specialized filters, and filters get increasingly specialized in deeper layers of the network.
When transferring the pointwise layers, we observed convergence issues across all datasets, even in selffer models, suggesting optimization difficulties. This may be attributed to the higher parameter count in pointwise layers, potential sparsity effects, and restricted permutation symmetries compared to depthwise layers. These findings align with prior work on fragile co-adaptation in neural networks (Yosinski et al. 2014), where freezing certain layers can create challenging loss landscapes for training the remaining parameters. Future work could further investigate the specific mechanisms behind these optimization challenges.